40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 13 17:37:43 2014, Jenne, Update, General, Game plan 2014_Aug_13.pdf
    Reply  Thu Aug 14 02:51:46 2014, Jenne, Update, General, Game plan Y_ALS_FINE_PHASE_OUT_noisy_13Aug2014.pdf
    Reply  Thu Aug 14 14:58:03 2014, Jenne, Update, General, Updated game plan 2014_Aug_14.pdf
       Reply  Thu Aug 14 15:42:29 2014, Koji, Update, General, Updated game plan 
          Reply  Thu Aug 14 15:51:37 2014, Jenne, Update, General, Updated game plan 
             Reply  Thu Aug 14 18:02:11 2014, Koji, Update, General, Updated game plan 
                Reply  Thu Aug 14 18:05:05 2014, Jenne, Update, General, Updated game plan 
    Reply  Fri Aug 15 20:26:37 2014, ericq, Update, General, Game plan 
Message ID: 10377     Entry time: Wed Aug 13 17:37:43 2014     Reply to this: 10382   10383   10404
Author: Jenne 
Type: Update 
Category: General 
Subject: Game plan 

2014_Aug_13.pdf

Here's the game plan for things that we need to do to get this IFO locked up. 

Red is for things that should be done today, or tomorrow if they don't get finished today (eg. laser mode hopping temperature check).  Orange is for things that will become red once the current red things are gone (eg. inferring the POP QPD gouy phase, and moving it to minimized PRM information).  Green is for things that we'd like to do, but aren't high priority (eg. X green mode matching).  Blue is for things that we should remember, but not plan on working on soon (eg. putting PZTs on the Yend table for green).

TODAY so far:

Q already did the tweak up of the PSL SHG crystal alignment.  HE SHOULD ELOG ABOUT THIS.  What was the final power of green that you got?  Do we have any record of a previous measurement to compare to?

Q helped me install PDA55s on each of the lasers (I did the ends, he did the PSL) so that we could do the mode hop temperature check.  For the Yend, I took the leakage transmission through the first Y1 steering mirror after the laser. This beam was dumped, so I replaced the dump with a PDA55. For the Xend, the equivalent mirrors are too close to the edge of the table, so I put in a spare Y1, and reflect most of the light to a beam dump.  The leakage transmission then goes to a PDA55.  Note that for both of these cases, no alignment of main laser path mirrors was touched, so we should just be able to remove them when we're through.  For the PSL, I believe that Q took the rejected light from one of the PBSes before the PMC.  He mentioned that he bumped something, so had to realign the beam into the PMC, but that he was able to get the transmission back up to 0.802, when we were seeing it in the mid 0.7's for the last several days.

The end temporary PDs are using the TRX / TRY cables, so we will be looking at the C1:LSC-TR[x,y] channels for the power of the end lasers.  The PSL's temporary PD is connected to the PMC REFL cable.  For the end PDs, since I had filter banks available, I shuttered the end lasers and removed the dark offset.  I then changed the gains to 1, so the values are in raw counts.  The usual transmission normalization gains are noted in one of the control room notebooks.

I did a slow ezcastep and ramped the temperature of all 3 lasers over about an hour.  I'll write a separate elog about how that went.

ELOG V3.1.3-